The Torrey Pine Corporation’s purchases from suppliers in a quarter are equal to 75 percent of the next quarter’s forecast sales. The payables period is 60 days. Wages, taxes, and other expenses are 20 percent of sales, and interest and dividends are $90 per quarter. No capital expenditures are planned.

Projected quarterly sales are shown here:

  Q1   Q2   Q3   Q4
  Sales $ 2,100 $ 2,400 $ 2,100 $ 1,800

Sales for the first quarter of the following year are projected at $2,430. Calculate the company’s cash outlays by completing the following (Do not round intermediate calculations. Round your answers to 2 decimal places, e.g., 32.16.):

Q1 Q2 Q3 Q4
  Payment of accounts $    $    $    $   
  Wages, taxes, other expenses            
  Long-term financing expenses
     (interest and dividends)
           
      Total $
Answer #1

Answer

  • Payment to supplier is 60 days means that payment for one quarter is made within that quarter only.

Q1

Q2

Q3

Q4

  Payment of accounts

$1,800

$1,575

$1,350

$1,823

  Wages, taxes, other expenses

$420

$480

$420

$360

  Long-term financing expenses

$90

$90

$90

$90

     (interest and dividends)

      Total

$2,310

$2,145

$1,860

$2,273

--Working

Q1

Q2

Q3

Q4

  Payment of accounts

=2400*75%

=2100*75%

=1800*75%

=2430*75%

  Wages, taxes, other expenses

=2100*20%

=2400*20%

=2100*20%

=1800*20%

  Long-term financing expenses

90

90

90

90

     (interest and dividends)